Problem 5A-7 Cost Behavior; High-Low Method; Contribution Format Income Statement (LO5-10] Morrisey & Brown, Ltd., of SydneyRequired: 1. By analyzing the data from the companys income statements, classify each of its expenses (including cost of go
Answer #1
1) Expenses Classification
cost of goods sold Variable
Advertising expense fixed
shipping expense mixed
Salaries and commission mixed
insurance expense fixed
Depreciation expense fixed
2) High low method
shipping expense
units cost
high 8,750 99,500
low 7,750 89,500
difference 1,000 10,000
variable expense per unit 10000/1000 = $10
fixed expense =99500-8750*10= 12000
Salaries and commission
units cost
high 8,750 155,000
low 7,750 140,000
difference 1,000 15,000
Variable expense per unit = 15000/1000=$15
fixed expense = 155000-8750*15= 23750
cost formula
Variable cost Fixed cost                    Formula                                                       
shipping expense 10 per unit 12,000 Y= 12,000 + 10 per unit
Salaries & commission 15 per unit 23,750 Y= 23,750 + 15 per unit
3) Sales in units 8750
Sales 875,000
Variable expense
Cost of goods sold 525,000
Shipping expense 87,500
Salaries and commission 131250
743,750
Contribution margin 131,250
Fixed expense
Advertising expense 28,200
Shipping expense 12,000
Salaries & commission 23,750
insurance expense 7,200
Depreciation expense 21,600
92,750
Net income 38,500